McEwen Company has budgeted the following credit sales during the last four months of the​ year: September,​ $14,000; October,​ $20,000; November​ $16,000; December,​ $33,000. Experience has shown that payment for the credit sales is received as​ follows: 30% in the month of​ sale, 55% in the first month after​ sale, 10% in the second month after​ sale, and​ 5% uncollectible. How much cash can McEwen Company expect to collect in November as a result of credit​ sales?

A.

​$18,850

B.

​$17,200

C.

​$15,800

D.

​$12,400
Answer #1

Cash Realisable on september sale in november = $14000 * 10% = $1,400

Cash Realisable on october sale in november = $ 20000 * 55% = $11,000

Cash Realisable on november sale in november = $16000 * 30% = $4,800

Total cash realisable in november as a result of credit sales = $1,400 + $11,000 + $4,800 = $17,200

So answer is option B = $17,200